Which system of equations cannot be
directly solved by applying the elimination
method?
A. 12𝑥 − 11𝑦 = −4
−6𝑥 − 11𝑦 = 7
B. 𝑥 − 11𝑦 = −6
𝑥 + 12𝑦 = −7
C. 7𝑥 + 11𝑦 = 12
−6𝑥 − 4𝑦 = −21
D. 𝑥 + 11𝑦 = 7
11𝑦 − 12𝑥 = 4

Found 3 solutions by Fombitz, advanced_Learner, MathTherapy:
Answer by Fombitz(32388)

Look at the x and y coefficients for each system.
The ones that have equal x or y coefficients can be solved directly.
Example: For A, both equations have -11y.
So this system can be solved directly.
Carefully look through the remaining three and find the one which doesn't have matching x or y coefficients.

Answer by advanced_Learner(501)

B AND D
what do you get if you add the constant term.

Answer by MathTherapy(10552)

Which system of equations cannot be
directly solved by applying the elimination
method?
A. 12𝑥 − 11𝑦 = −4
−6𝑥 − 11𝑦 = 7
B. 𝑥 − 11𝑦 = −6
𝑥 + 12𝑦 = −7
C. 7𝑥 + 11𝑦 = 12
−6𝑥 − 4𝑦 = −21
D. 𝑥 + 11𝑦 = 7
11𝑦 − 12𝑥 = 4
It's NOT B and D. It's C, because A, B, and D CAN be solved, DIRECTLY, using the elimination method.